How do I solve for x in this problem?
1/x + 1/(x - 3) = 2/(x + 3)

Answer by checkley77(12844)

1/x + 1/(x - 3) = 2/(x + 3)
1/X+1/(X-3)-2/(X+3)=0
[(X-3)(X+3)+X(X+3)-2X(X-3)]/X(X-3)(X+3)=0
[X^2-9+X^2+3X-2X^2+6X]/X(X-3)(X+3)=0
[9X-9]/X(X-3)(X+3)=0
9[X-1]/X(X-3)(X+3)=0